# A Robust and Non-Iterative Tensor Decomposition Method with Automatic Thresholding

### v2

@@ Abstract
Recent advances in IoT and biometric sensing technologies have led to the generation of massive and high-dimensional tensor data, yet achieving accurate and efficient low-rank approximation remains a major challenge. Existing tensor decomposition methods typically require prior specification of the tensor rank and rely on iterative optimization, which often results in heavy computational costs and dependence on the analyst’s expertise. In this study, we propose a novel low-rank approximation method for tensor data that requires neither prior rank specification nor iterative optimization. The proposed method performs statistical singular value hard thresholding on the mode-wise unfolded matrices to automatically extract only statistically significant components, thereby achieving noise reduction while preserving the intrinsic tensor structure. Theoretically, the optimal threshold for each mode is derived based on the asymptotic properties of the Marčenko-Pastur distribution. Simulation experiments demonstrate that the proposed method outperforms conventional approaches such as Higher-Order Singular Value Decomposition, Higher-Order Orthogonal Iteration, and Tucker-L2E in terms of both estimation accuracy and computational efficiency. These results indicate that our method provides an effective and theoretically grounded framework for automatic, non-iterative, and analyst-independent tensor decomposition.
@@
Multivariate statistics , Machine learning in OR , Data science , Robust optimization , tensor
@@ 1 Introduction
In modern society, vast amounts of data are continuously collected from various information sources such as sensors, social networks, and biometric information systems. These data often have tensor structures that simultaneously include temporal, spatial, and categorical relationships, serving as an important foundation for capturing phenomena from multiple perspectives (Kolda and Bader,, 2009). However, there are constraints such as measurement errors and environmental fluctuations in real-world data collection environments, and the acquired data inherently contain noise. This noise obscures meaningful patterns and becomes a factor that undermines the reliability of analytical results (Haven et al.,, 2012).
@@
In recent years, in the field of Operations Research (OR), the importance of analysis utilizing real-world data has been further increasing. However, due to the limitations of measurement accuracy and environmental conditions, noise is inevitable in real-world data, and appropriately handling its influence has become an important issue (Haven et al.,, 2012). Previous studies have proposed various methods to reduce the effects of noise, such as the evaluation of feature selection for improving classification accuracy (Piramuthu,, 2004), the proposal of dynamic resampling strategies in evolutionary optimization (Syberfeldt et al.,, 2010), the improvement of price forecasting accuracy through wavelet transforms (Haven et al.,, 2012), and the construction of models capable of estimating both short-term and long-term inefficiencies (Badunenko and Kumbhakar,, 2016).
@@
All of these studies aim to suppress the influence of noise and to achieve more accurate and reliable analyses; however, many of them have been discussed within matrix-based frameworks. In modern data environments, however, high-dimensional data are becoming common, as exemplified by sensor analysis and biometric signal processing, where diverse elements such as time, space, and category are intertwined. In such cases, it is difficult for matrix-based analytical methods to avoid the loss of structural information. In particular, as the number of variables and observation points increases, the so-called “curse of dimensionality” becomes apparent, leading to an increase in computational burden and the risk of overlooking latent structures (Kilmer et al.,, 2019; Auddy et al.,, 2024).
@@
To address these challenges, tensor analysis, which enables analysis while preserving the multidimensional structure of data, has been attracting increasing attention. Because tensors can handle high-dimensional data while maintaining relationships among different modes, they are particularly effective in applications that simultaneously consider multiple dimensions such as time, space, and category (e.g., demand forecasting and anomaly detection) (Kolda and Bader,, 2009). Furthermore, it has been shown that low-rank approximation can suppress the influence of noise and missing data, thereby achieving more accurate predictions and robust anomaly detection (Acar et al.,, 2011; Zhang and Aeron,, 2016). Against this background, a promising strategy has been positioned as one that separates observed data into signal and noise components and removes noise by estimating the signal. In particular, tensor decomposition methods such as CP decomposition, Tucker decomposition, and Tensor-Train decomposition have been attracting attention as powerful tools for noise removal and structural recovery in multidimensional data (Carroll and Chang,, 1970; Harshman,, 1970; Tucker,, 1966; Oseledets,, 2011).
@@ 3.1 Computational Cost Analysis
This section presents a theoretical analysis of the computational complexity of the proposed TARST algorithm and derives its WCTR. Following standard asymptotic notation, $\mathcal{O}(\cdot)$ and $\Omega(\cdot)$ denote the upper and lower bounds of computational complexity, respectively. Formally, for two positive functions $f(n)$ and $g(n)$:
@@
The evaluation focuses on three key aspects. First, following the worst-case complexity framework proposed by Absi and Heuvel, (2019), the theoretical upper bound of TARST’s computational cost is derived by decomposing and summing its mode-wise operations, thereby establishing its asymptotic order. Second, a closed-form expression is formulated under isotropic conditions, where all mode sizes are equal, providing a simplified yet generalizable representation that clarifies how the bound scales with tensor dimension and rank ratio. Third, the analytical conditions guaranteeing the existence of a finite and polynomial computational bound are investigated to ensure the theoretical soundness and stability of the proposed method. The results section summarizes these analyses, offering a comprehensive theoretical understanding of TARST’s computational characteristics and upper-bound behavior.
@@ 3.2 Reconstruction Accuracy Analysis
The objective of this analysis is to quantitatively evaluate the denoising and reconstruction performance of the proposed TARST method in comparison with existing tensor decomposition approaches. Using synthetic datasets with known ground-truth structures allows direct assessment of how accurately each method reconstructs the underlying signal under different noise and outlier conditions.
All experiments were conducted in MATLAB R2025b using Tensor Toolbox v3.6 and the official TuckerL2E-master implementation available on GitHub. Default Tucker-L2E parameters were used, and the random seed was fixed (rng(0)) to ensure reproducibility.
The true tensors were designed to have predefined low-rank structures, which served as the ground-truth ranks in the experiments. For a fair comparison, rank-specified methods (HOSVD, HOOI) were provided with the correct rank information corresponding to the true low-rank structure, while the proposed TARST automatically estimated the rank for each mode without prior specification.
@@
Five methods were compared in total: Baseline, HOSVD, HOOI, Tucker-L2E, and the proposed TARST. The Baseline simply uses the noisy observation tensor as a reference for the effect of noise. HOSVD (De Lathauwer et al., 2000a, ) performs Tucker decomposition with a fixed rank in a non-iterative manner, while HOOI (De Lathauwer et al., 2000b, ) refines the HOSVD result iteratively through Alternating Least Squares (ALS) optimization. Tucker-L2E (Chen et al.,, 2020) is a robust extension of the Tucker model based on the $L_{2}E$ estimator, designed to mitigate the influence of outliers. Finally, the proposed TARST performs automatic rank estimation and singular value hard thresholding for each mode without iterative optimization, enabling fast and parameter-free tensor reconstruction.
Two types of synthetic experiments were conducted to evaluate performance under different noise and outlier conditions. In the first experiment (Pattern 1), Gaussian noise was added to assess denoising capability. Two tensor sizes were used: $10\times 10\times 10$ and $50\times 50\times 50$. Each tensor was generated from a normal distribution with mean 10 and standard deviations of either 2 or 0.25, resulting in four total combinations. Gaussian noise with standard deviation $\sigma\in[10^{-1},10^{1}]$ (log scale) was added as
$\mathcal{Y}=\mathcal{X}+\sigma\mathcal{E},$
@@
where $\mathcal{Y}$ is the observed tensor, $\mathcal{X}$ the ground truth, and $\mathcal{E}\sim\mathcal{N}(0,1)$ the Gaussian noise. While $\sigma$ is known in these synthetic experiments to control the noise level, the proposed TARST does not require prior knowledge of $\sigma$ during decomposition, as the threshold is automatically determined from the empirical singular-value distribution. A smaller $\sigma$ corresponds to low-noise conditions, whereas a larger $\sigma$ represents high-noise conditions. Each condition was repeated five times, and the mean relative reconstruction squared error(RRSE) with 95% confidence intervals was reported.
@@
Reconstruction accuracy was evaluated using the RRSE(Zhao et al.,, 2015):
$\mathrm{RRSE}(\hat{\mathcal{X}},\mathcal{X})=\frac{\|\hat{\mathcal{X}}-\mathcal{X}\|_{F}}{\|\mathcal{X}\|_{F}}.$
A smaller RRSE indicates a more accurate recovery of the true underlying tensor structure.
@@ 4.1.1 Derivation of the Theoretical Upper Bound
The computational cost of the proposed TARST algorithm can be decomposed into three major phases for each tensor mode: (1) SVD of the mode-unfolded matrices, (2) Hard thresholding of singular values, and (3) Tucker-type reconstruction using the post-threshold components. Given a tensor with $N$ modes, the overall computational cost can be expressed as
@@
By aggregating these three components, the overall theoretical upper bound of TARST can be expressed as
@@
$T_{\mathrm{TARST}}=\mathcal{O}\!\Bigg(\sum_{k=1}^{N}\Bigg[\min\left\{I_{k}^{2}I_{(-k)},I_{k}I_{(-k)}^{2}\right\}+\min\left\{I_{k},I_{(-k)}\right\}+\frac{P\,r_{(-k)}}{I_{k}}\Bigg]\Bigg).$
@@
4.1.2 Closed-Form Expression under Isotropic Conditions
The general form of the computational upper bound derived in the previous subsection can be further simplified under an isotropic condition, in which all tensor modes are assumed to have equal dimensions. Specifically, let the mode sizes and post-threshold ranks be defined as $I_{k}=I$ and $r_{k}=r$ for all $k=1,\dots,N$. Under this assumption, the total number of tensor elements is $P=I^{N}$, and each mode-unfolded matrix has the size $I\times I^{N-1}$.
By substituting these conditions into the general expression of the upper bound, the theoretical computational complexity of TARST can be expressed as
$T_{\mathrm{TARST}}=\mathcal{O}\!\Bigg(N\bigg[I^{N+1}+I+I^{N-1}r^{N-1}\bigg]\Bigg).$
In this formulation, the first term $I^{N+1}$ corresponds to the cost of the SVD for each mode, the second term $I$ represents the computational cost of the hard-thresholding step, and the third term $I^{N-1}r^{N-1}$ corresponds to the reconstruction phase, which depends on the post-threshold ranks. This closed-form representation clearly characterizes the asymptotic behavior of the TARST computational cost as a function of the tensor dimension $I$, the number of modes $N$, and the rank parameter $r$.
To further clarify the scaling relationship, a rank ratio $\gamma=r/I$ can be introduced, yielding the following expression:
$T_{\mathrm{TARST}}=\mathcal{O}\!\Bigg(NI^{N+1}\bigg[1+I^{-N}+\gamma^{N-1}\bigg]\Bigg).$
This closed-form expression provides an analytical foundation for interpreting the dependency of TARST’s computational cost on tensor size and rank ratio, and serves as the theoretical basis for the subsequent analysis of the tightness visualization and the existence conditions of the computational upper bound.
4.1.3 Investigation of Bound Existence Conditions
In this subsection, we theoretically investigate the conditions under which the computational upper bound of TARST exists and remains finite. Based on the previous derivations, the overall computational cost of TARST can be expressed as
$T_{\mathrm{TARST}}=\mathcal{O}\!\Bigg(\sum_{k=1}^{N}\Bigg[\min\left\{I_{k}^{2}I_{(-k)},I_{k}I_{(-k)}^{2}\right\}+\min\left\{I_{k},I_{(-k)}\right\}+\frac{P\,r_{(-k)}}{I_{k}}\Bigg]\Bigg),$
where $I_{(-k)}=\prod_{j\neq k}I_{j}$, $r_{(-k)}=\prod_{j\neq k}r_{j}$, and $P=\prod_{k=1}^{N}I_{k}$.
Proposition 1 (Finite upper bound of each computational phase). Each component of the TARST cost function admits a finite upper bound under the following conditions:
  1. (i)
    (SVD phase) The term $\min\left\{I_{k}^{2}I_{(-k)},I_{k}I_{(-k)}^{2}\right\}$ is always finite for bounded mode dimensions $I_{k}<\infty$, and is polynomial in $I_{k}$ when the number of modes $N$ is fixed.
  2. (ii)
    (Hard-thresholding phase) The term $\min\left\{I_{k},I_{(-k)}\right\}$ is a linear function of mode size and remains finite by definition.
  3. (iii)
    (Reconstruction phase) The term $\frac{P\,r_{(-k)}}{I_{k}}$ remains finite if the post-threshold rank satisfies $1\leq r_{k}<I_{k}$, which holds by the construction of the thresholding process.
Proposition 2 (Polynomial-time condition). Let the post-threshold ranks satisfy $r_{k}=\mathcal{O}\left(I_{k}^{\alpha_{k}}\right)$ with $0\leq\alpha_{k}<1$ for each mode $k$. Then the total computational cost of TARST satisfies
$T_{\mathrm{TARST}}=\mathcal{O}\!\Bigg(\sum_{k=1}^{N}I_{k}^{N+1}+\sum_{k=1}^{N}I_{k}+\sum_{k=1}^{N}P\,I_{k}^{\alpha_{(-k)}-1}\Bigg),\quad\text{where }\alpha_{(-k)}=\sum_{j\neq k}\alpha_{j}.$
@@ 4.1.3 Investigation of Bound Existence Conditions
Since $\alpha_{k}<1$ for all $k$, the term $I_{k}^{\alpha_{(-k)}-1}$ is bounded, and the total cost satisfies $T_{\mathrm{TARST}}=\mathcal{O}\left(P^{1+\varepsilon}\right)$ with $\varepsilon<1$. Hence, TARST operates within polynomial time with respect to the input size $P$.
Proposition 3 (Boundedness of the Worst-Case Time Ratio). Let the WCTR be defined as
$\mathrm{WCTR}(\mathrm{TARST})=\sup_{\mathcal{Y}}\frac{T_{\mathrm{TARST}}(\mathcal{Y})}{T_{\mathrm{opt}}(\mathcal{Y})},\qquad T_{\mathrm{opt}}(\mathcal{Y})=\Omega(P).$
If $T_{\mathrm{TARST}}=\mathcal{O}\left(P^{1+\varepsilon}\right)$ with $\varepsilon<1$, then
$\mathrm{WCTR}(\mathrm{TARST})=\mathcal{O}(P^{\varepsilon}),$
which guarantees that the ratio is finite for any bounded $\varepsilon<1$. Therefore, the theoretical upper bound of TARST exists and remains stable across all feasible tensor configurations.
Corollary 1 (Sufficient conditions for bounded and polynomial complexity). The computational cost of TARST admits a finite and polynomial upper bound under the following sufficient conditions:
$\begin{cases}N<\infty,\\ 1\leq r_{k}<I_{k},\\ r_{k}=\mathcal{O}\left(I_{k}^{\alpha_{k}}\right),\ \ 0\leq\alpha_{k}<1.\end{cases}$
Under these assumptions, it follows that
$T_{\mathrm{TARST}}=\mathcal{O}\left(P^{1+\varepsilon}\right),\qquad\varepsilon<1,$
demonstrating that TARST possesses a stable computational complexity bounded by a finite polynomial order with respect to the input size.

### v3

@@ Abstract
Recent progress in IoT and biometric sensing technologies has produced massive, high-dimensional tensor data, but achieving accurate and efficient low-rank approximation remains challenging. Most existing tensor decomposition methods require predefined ranks and iterative optimization, leading to high computational costs and reliance on analyst expertise. This study proposes a new tensor low-rank approximation method that eliminates both prior rank specification and iterative optimization. The method applies statistical singular value hard thresholding to each mode-wise unfolded matrix to automatically extract statistically significant components, reducing noise while preserving intrinsic structure. Theoretically, optimal thresholds for each mode are derived from the asymptotic properties of the Marčenko–Pastur distribution. Simulations show that the proposed method outperforms conventional approaches—HOSVD, HOOI, and Tucker-L2E—in both estimation accuracy and computational efficiency. These results indicate that the method offers a theoretically grounded, fully automatic, and non-iterative framework for tensor decomposition.
@@
Multivariate statistics , Tensor decomposition , Automatic thresholding , Noise reduction , Data science
@@ 1 Introduction
Vast amounts of data are continuously collected from diverse sources, including sensors, social networks, and biometric systems. Such data often exhibit tensor structures capturing temporal, spatial, and categorical relationships, forming a foundation for multi-perspective analysis(Kolda and Bader,, 2009). However, real-world data collection involves measurement errors and environmental fluctuations, inherently introducing noise. Such noise obscures meaningful patterns and undermines analytical reliability (Haven et al.,, 2012).
@@
In Operations Research (OR), the importance of real-world data analysis has been steadily increasing. Yet, noise remains inevitable due to measurement and environmental limitations, making its proper treatment a key issue(Haven et al.,, 2012). Prior studies have proposed various approaches to mitigate noise effects, including feature selection for improved classification accuracy (Piramuthu,, 2004), the proposal of dynamic resampling strategies in evolutionary optimization (Syberfeldt et al.,, 2010), the improvement of price forecasting accuracy through wavelet transforms (Haven et al.,, 2012), and the construction of models capable of estimating both short-term and long-term inefficiencies (Badunenko and Kumbhakar,, 2016).
@@
While all these studies sought to suppress noise and improve accuracy, most relied on matrix-based frameworks. However, modern data environments increasingly involve high-dimensional data, such as sensor or biometric signals that intertwine temporal, spatial, and categorical elements. In such cases, it is difficult for matrix-based analytical methods to avoid the loss of structural information. In particular, as the number of variables and observation points increases, the so-called “curse of dimensionality” becomes apparent, leading to an increase in computational burden and the risk of overlooking latent structures (Kilmer et al.,, 2019; Auddy et al.,, 2024).
@@
To address these challenges, tensor analysis has gained attention for its ability to preserve multidimensional data structures during analysis. Because tensors can handle high-dimensional data while maintaining relationships among different modes, they are particularly effective in applications that simultaneously consider multiple dimensions such as time, space, and category (e.g., demand forecasting and anomaly detection) (Kolda and Bader,, 2009). Furthermore, it has been shown that low-rank approximation can suppress the influence of noise and missing data, thereby achieving more accurate predictions and robust anomaly detection (Acar et al.,, 2011; Zhang and Aeron,, 2016). Against this background, a promising strategy has been positioned as one that separates observed data into signal and noise components and removes noise by estimating the signal. In particular, tensor decomposition methods such as CP decomposition, Tucker decomposition, and Tensor-Train decomposition have been attracting attention as powerful tools for noise removal and structural recovery in multidimensional data (Carroll and Chang,, 1970; Harshman,, 1970; Tucker,, 1966; Oseledets,, 2011).
@@ 3.1 Computational Cost Analysis
This section presents a theoretical analysis of the computational complexity of the proposed TARST algorithm and derives its Worst-Case Time Ratio (WCTR).
Following standard asymptotic notation, $\mathcal{O}(\cdot)$ and $\Omega(\cdot)$ denote the upper and lower bounds of computational complexity, respectively. Formally, for two positive functions $f(n)$ and $g(n)$:
@@
The evaluation focuses on the derivation of the theoretical upper bound and the investigation of the existence conditions for a finite polynomial bound of TARST. Specifically, the theoretical upper bound of TARST’s computational cost is derived by decomposing and summing its mode-wise operations under the worst-case complexity framework Absi and Heuvel, (2019), thereby establishing its asymptotic order. Furthermore, the analytical conditions under which this bound remains finite and polynomial are examined to ensure the theoretical soundness and computational stability of the proposed method. The results section summarizes these analyses, providing a comprehensive understanding of TARST’s computational characteristics and upper-bound behavior.
@@ 3.2 Reconstruction Accuracy Analysis
The objective of this analysis is to quantitatively evaluate the denoising and reconstruction performance of the proposed TARST method in comparison with existing tensor decomposition approaches. Using synthetic datasets with known ground-truth structures allows direct assessment of how accurately each method reconstructs the underlying signal under different noise and outlier conditions. All experiments were conducted in MATLAB R2025b using Tensor Toolbox v3.6 and the official TuckerL2E-master implementation available on GitHub. Default Tucker-L2E parameters were used, and the random seed was fixed (rng(0)) to ensure reproducibility.
@@
The true tensors were designed to have predefined low-rank structures, which served as the ground-truth ranks in the experiments. For a fair comparison, rank-specified methods (HOSVD, HOOI) were provided with the correct rank information corresponding to the true low-rank structure, while the proposed TARST automatically estimated the rank for each mode without prior specification. Five methods were compared in total: Baseline, HOSVD, HOOI, Tucker-L2E, and the proposed TARST. The Baseline simply uses the noisy observation tensor as a reference for the effect of noise. HOSVD (De Lathauwer et al., 2000a, ) performs Tucker decomposition with a fixed rank in a non-iterative manner, while HOOI (De Lathauwer et al., 2000b, ) refines the HOSVD result iteratively through Alternating Least Squares (ALS) optimization. Tucker-L2E (Chen et al.,, 2020) is a robust extension of the Tucker model based on the $L_{2}E$ estimator, designed to mitigate the influence of outliers. Finally, the proposed TARST performs automatic rank estimation and singular value hard thresholding for each mode without iterative optimization, enabling fast and parameter-free tensor reconstruction.
@@
Two types of synthetic experiments were conducted to evaluate performance under different noise and outlier conditions. In the first experiment (Pattern 1), Gaussian noise was added to assess denoising capability. Two tensor sizes were used: $10\times 10\times 10$ and $50\times 50\times 50$. Each tensor was generated from a normal distribution with mean 10 and standard deviations of either 2 or 0.25, resulting in four total combinations. Gaussian noise with standard deviation $\sigma\in[10^{-1},10^{1}]$ (log scale) was added as $\mathcal{Y}=\mathcal{X}+\sigma\mathcal{E}$, where $\mathcal{Y}$ is the observed tensor, $\mathcal{X}$ is the ground truth, and $\mathcal{E}\sim\mathcal{N}(0,1)$ is Gaussian noise. While $\sigma$ is known in these synthetic experiments to control the noise level, the proposed TARST does not require prior knowledge of $\sigma$ during decomposition, as the threshold is automatically determined from the empirical singular-value distribution. A smaller $\sigma$ corresponds to low-noise conditions, whereas a larger $\sigma$ represents high-noise conditions. Each condition was repeated five times, and the mean relative reconstruction squared error (RRSE) with 95% confidence intervals was reported. The RRSE (Zhao et al.,, 2015) is defined as
$\mathrm{RRSE}(\hat{\mathcal{X}},\mathcal{X})=\frac{\|\hat{\mathcal{X}}-\mathcal{X}\|_{F}}{\|\mathcal{X}\|_{F}},$
where $\hat{\mathcal{X}}$ denotes the reconstructed tensor and $\mathcal{X}$ the ground truth. A smaller RRSE indicates a more accurate recovery of the true underlying tensor structure.
@@ 4.1.1 Derivation of the Theoretical Upper Bound
The computational cost of the proposed TARST algorithm can be divided into three major phases for each tensor mode: (1) SVD of the mode-unfolded matrices, (2) hard thresholding of singular values, and (3) Tucker-type reconstruction using the post-threshold components. For a tensor with $N$ modes, the overall computational cost is expressed as
@@
By aggregating these three components, the overall theoretical upper bound of TARST is given by
@@
$T_{\mathrm{TARST}}=\mathcal{O}\!\Bigg(\sum_{k=1}^{N}\Bigg[\min\left\{I_{k}^{2}I_{(-k)},\,I_{k}I_{(-k)}^{2}\right\}+\min\left\{I_{k},\,I_{(-k)}\right\}+\frac{P\,r_{(-k)}}{I_{k}}\Bigg]\Bigg).$ (1)
@@
4.1.2 Polynomial-Time Bound under General Conditions
We next derive the theoretical upper bound of TARST under general, anisotropic conditions without assuming equal tensor mode sizes. Let the tensor have $N<\infty$ modes, and denote its mode dimensions and post-threshold ranks as $I_{k}$ and $r_{k}$, respectively. Assume that the ranks increase more slowly than their corresponding mode dimensions, i.e.,
$r_{k}=\mathcal{O}(I_{k}^{\alpha_{k}}),\qquad 0\leq\alpha_{k}<1,\quad k=1,2,\dots,N.$
Let $P=\prod_{k=1}^{N}I_{k}$ denote the total number of tensor elements. Then, the following proposition holds.
Proposition 4.1 (Polynomial-Time Upper Bound under General Anisotropic Conditions).
Under the above assumptions, the total computational cost of TARST satisfies
$T_{\mathrm{TARST}}=\mathcal{O}(P^{1+\varepsilon}),$
@@ 4.1.2 Polynomial-Time Bound under General Conditions
where $\varepsilon=\max\{1/2,\,\alpha_{\max}\}$ and $\alpha_{\max}=\max_{k}(\alpha_{k})<1$. Hence, TARST operates within polynomial time with respect to the input size $P$.
Proof. The computational upper bound in Eq. (1) consists of three major terms: the SVD term, the hard-thresholding term, and the reconstruction term. Each term is now analyzed separately.
(1) SVD term. The asymptotic cost of the SVD for each mode is
$T_{\mathrm{SVD},k}=\mathcal{O}\!\left(\min\left\{I_{k}^{2}I_{(-k)},\,I_{k}I_{(-k)}^{2}\right\}\right).$
Because $P=\prod_{j=1}^{N}I_{j}=I_{k}I_{(-k)}$, this expression can be rewritten in terms of $I_{k}$ and $P$ as
$\min\left\{I_{k}P,\,\frac{P^{2}}{I_{k}}\right\}.$
By applying the arithmetic–geometric mean inequality, which holds since all elements are positive, we can set a loose but valid upper bound as
$\min\left\{I_{k}P,\,\frac{P^{2}}{I_{k}}\right\}\leq\sqrt{(I_{k}P)\!\left(\frac{P^{2}}{I_{k}}\right)}=P^{3/2}.$
Thus, the SVD term is upper-bounded by $\mathcal{O}(P^{3/2})$.
(2) Hard-thresholding term. The thresholding phase has asymptotic cost
$T_{\mathrm{thresh},k}=\mathcal{O}\!\left(\min\{I_{k},I_{(-k)}\}\right)=\mathcal{O}\!\left(\min\left\{I_{k},\frac{P}{I_{k}}\right\}\right).$
Similarly, using the same bounding technique,
$\min\left\{I_{k},\frac{P}{I_{k}}\right\}\leq\sqrt{I_{k}\left(\frac{P}{I_{k}}\right)}=P^{1/2}.$
Hence, the hard-thresholding term contributes an order of $\mathcal{O}(P^{1/2})$, which is asymptotically smaller than the SVD term.
(3) Reconstruction term. The reconstruction phase is expressed as
$T_{\mathrm{recon},k}=\mathcal{O}\!\left(\frac{P\,r_{(-k)}}{I_{k}}\right).$
To evaluate $r_{(-k)}$, we assume $r_{j}=\mathcal{O}(I_{j}^{\alpha_{j}})$ for each mode, where $0\leq\alpha_{j}<1$. Let $\alpha_{\max}=\max_{j}(\alpha_{j})$. Then, for any $j$, we have $r_{j}\leq C_{j}I_{j}^{\alpha_{\max}}$, where $C_{j}$ is a constant depending only on mode $j$. Therefore,
$r_{(-k)}=\prod_{j\neq k}r_{j}\leq\left(\prod_{j\neq k}C_{j}\right)\left(\prod_{j\neq k}I_{j}^{\alpha_{\max}}\right)=C\cdot I_{(-k)}^{\alpha_{\max}},$
where $C=\prod_{j\neq k}C_{j}$ is a constant. Since $I_{(-k)}=P/I_{k}$, this becomes $r_{(-k)}\leq C(P/I_{k})^{\alpha_{\max}}$. Substituting this into the reconstruction term yields
$\frac{P\,r_{(-k)}}{I_{k}}\leq C\cdot\frac{P}{I_{k}}\left(\frac{P}{I_{k}}\right)^{\alpha_{\max}}=C\cdot P^{1+\alpha_{\max}}I_{k}^{-(1+\alpha_{\max})}.$
Because $I_{k}\geq 1$, the inverse power term $I_{k}^{-(1+\alpha_{\max})}$ is bounded above by one, resulting in
$T_{\mathrm{recon},k}=\mathcal{O}(P^{1+\alpha_{\max}}).$
(4) Summary. Combining the three terms, we obtain
$T_{\mathrm{TARST}}=\mathcal{O}\!\left(P^{3/2}+P^{1/2}+P^{1+\alpha_{\max}}\right).$
Among these, the dominant exponent is $\max\{3/2,1+\alpha_{\max}\}=1+\max\{1/2,\alpha_{\max}\}$. Letting $\varepsilon=\max\{1/2,\alpha_{\max}\}$, and noting that $\alpha_{\max}<1$, we conclude that
$T_{\mathrm{TARST}}=\mathcal{O}(P^{1+\varepsilon}),\qquad\varepsilon<1.$
Therefore, the theoretical computational complexity of TARST is bounded within polynomial time, and the exponent is strictly less than two. $\square$
Corollary 4.1.1 (Boundedness of the Worst-Case Time Ratio).
Let the optimal computational lower bound be $T_{\mathrm{opt}}=\Omega(P)$. Then, from the above proposition, the WCTR of TARST satisfies
$\mathrm{WCTR}(\mathrm{TARST})=\mathcal{O}(P^{\varepsilon}),\qquad\varepsilon<1.$
This result indicates that the WCTR of TARST always remains finite, and the computational growth rate is strictly linear with respect to the input size $P$. Therefore, TARST guarantees stable computational scalability within polynomial time, even under general anisotropic settings.
When the isotropic condition ($I_{k}=I$, $r_{k}=r$) is additionally assumed, the general expression derived above simplifies considerably. Because the constant multiplicative factor can be separated from the asymptotic order (and thus $N$ can be omitted), the computational cost reduces to
$T_{\mathrm{TARST}}=\mathcal{O}\!\Big(I^{N+1}+I+I^{N-1}r^{N-1}\Big).$
In this formulation, the first term $I^{N+1}$ corresponds to the SVD cost for each mode, the second term $I$ represents the hard-thresholding step, and the third term $I^{N-1}r^{N-1}$ denotes the reconstruction cost. If we further assume that $r=\mathcal{O}(I^{\alpha})$ with $0\leq\alpha<1$, the overall computational cost becomes
$T_{\mathrm{TARST}}=\mathcal{O}(P^{1+\frac{1}{N}}),$
which is consistent with the theoretical bound obtained under the general anisotropic setting. This isotropic formulation provides an intuitive interpretation of the asymptotic scaling behavior of TARST: the dominant term $I^{N+1}$ grows polynomially with respect to the tensor dimension, and the exponent $1+\tfrac{1}{N}$ becomes smaller as the tensor order $N$ increases, indicating that higher-order tensors enhance the relative computational stability of TARST.
4.1.3 Existence Conditions for the Polynomial-Time Bound
Based on the polynomial-time upper bound derived in Proposition 4.1, we now clarify the theoretical conditions under which a finite polynomial-time bound exists for TARST.
For a dense tensor $\mathcal{Y}\in\mathbb{R}^{I_{1}\times\cdots\times I_{N}}$ with $P=\prod_{k=1}^{N}I_{k}$ elements, let $r_{k}=\mathcal{O}(I_{k}^{\alpha_{k}})$ denote the post-threshold rank in each mode. Then, the computational cost of TARST satisfies
$T_{\mathrm{TARST}}=\mathcal{O}(P^{1+\varepsilon}),\qquad\varepsilon<1,$
provided that the following two general conditions hold:
  • (C1)
    The rank-growth exponents satisfy $\alpha_{k}<1$ for all $k=1,2,\dots,N$, ensuring that each mode-wise rank increases more slowly than its corresponding mode dimension. This prevents the reconstruction term from dominating the asymptotic order.
  • (C2)
    The tensor order $N$ is finite ($N<\infty$), so that the mode-wise summation in Eq. (1) contributes only a constant multiplicative factor to the total cost.
Under these assumptions, the exponent $\varepsilon=\max\{1/2,\alpha_{\max}\}$ with $\alpha_{\max}=\max_{k}(\alpha_{k})$ remains strictly less than one. Consequently, TARST is guaranteed to operate within polynomial time, and its computational complexity exponent is strictly smaller than two. This ensures that the algorithm achieves stable computational scalability even for high-order or anisotropic tensor data.